OpenStudy (openstudier):

A certain type of wrapping paper is sold in rolls of radius 5cm and has a cardboard core of radius 1 cm. If the paper is used until the thickness of the paper layer is half its original thickness, what fractional part of the original amount of paper is still on the roll?

OpenStudy (anonymous):

 then you need to find the area before and after pi * r^2 pi*5^2-pi*1^2=76.4 pi*3^2-pi*1^2=26.13 26.13/76.4=0.34 34%

OpenStudy (anonymous):

does that make sense?

OpenStudy (openstudier):

No. How do you figure out the second part?

OpenStudy (anonymous):

well i found the area of the whole part of the circle and then subtracted that from the cardboard part. This gives you just the area of the paper

OpenStudy (openstudier):

I understand that. What about the next part?

OpenStudy (anonymous):

please be more specific

OpenStudy (openstudier):

pi*3^2-pi*1^2=26.13

OpenStudy (anonymous):

"the paper is used until the thickness of the paper layer is half its original thickness" original thickness: 4 new thickness:2 new thickness + inner cardboar:3

OpenStudy (openstudier):

Why do you include the inner cardboard?

OpenStudy (anonymous):